Acetone is the main ingredient of nail polish remover. A solution is made up by adding 35.0 cm^3 of acetone (0.790 g cm^-3) to 50.0 cm^3 of ethyl alcohol (0.789 g cm^-3).Assuming volume are addictive, calculate,

  1. The mass percent of acetone in the solution
  2. The volume percent of ethyl alcohol in the solution
  3. The mole fraction of the acetone in the solution
1
Expert's answer
2020-08-30T02:22:46-0400

1.% mass of acetone is mass of solute split mass of solute=mass of solvent

multiply 100

"\frac{27.65}{27.65+39.45}\times100=\frac{27.65}{67.1}\times100=41.21" %


2.% volume of ethyl alcohol is volume of ethyl alcohol split total volume multiply 100

"\frac{50}{35+50}\times100=\frac{50}{85}\times100=58.82" %


3.mole fraction acetone is moles acetone split total moles)

"\frac{27.65}{58.08}=0.4761"


"\frac{39.45}{46.069}=0.8563"


"\frac{0.4761}{0.4761+0.8563}=\frac{0.4761}{1.33}=0.358"